OpenStudy (anonymous):

(x-68)^(2/5) = 81. There are apparently multiple answers, and I've only found one (59117.) Please help!

OpenStudy (anonymous):

Hello there.

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

-59117?

OpenStudy (anonymous):

Afraid I've already tried that one. I guess I'll give it one more try just for luck.

OpenStudy (anonymous):

well maybe not - it is a fractional root

OpenStudy (anonymous):

Nope nope nope.

OpenStudy (anonymous):

Yeah, I've no idea how I'm supposed to find the second answer.

OpenStudy (anonymous):

Give up? lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Well...I guess I'll just ask my bf when he comes online. Thank you though!

OpenStudy (anonymous):

okay, i found a way to graph it online and it doesn't seem like there would be another answer

OpenStudy (anonymous):

Doesn't surprise me. I really should've just gone to a tutoring session.

OpenStudy (anonymous):

yeah, that can help

OpenStudy (anonymous):

Anyway...not to be clingly/overly demanding...but...care to help me with another problem?

OpenStudy (anonymous):

Namely this: x^{3/8} - 12 x^{3/16} + 36 = 0

OpenStudy (anonymous):

clingy? it's math! not a problem, um... that's... a problem :-P

OpenStudy (anonymous):

I mean...I'm presumably supposed to use substitution.

OpenStudy (anonymous):

So say I replace x^(3/8) with y.

OpenStudy (anonymous):

u - 12u^(1/2) + 36?

OpenStudy (anonymous):

That gives me y - 12y^2 + 36, which is alot less intimidating

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Or wait...no. Is it 12y^(1/2)?

OpenStudy (anonymous):

y ^2 would back substitute to x ^(3/4)

OpenStudy (anonymous):

Oh...maybe that's why then...

OpenStudy (anonymous):

Wait...but 3/16 * 2 = 3/8.

OpenStudy (anonymous):

not if you replace y = x ^(3/8)

OpenStudy (anonymous):

But that wouldn't make sense. Because the object is to substitute another value in for the variable/exponent combination in the middle (i.e. 12x^3/16)

OpenStudy (anonymous):

x^{3/8} - 12 x^{3/16} + 36 = 0 would go to: y - 12y^2 + 36 = 0

OpenStudy (anonymous):

But 3/8 * 2 doesn't equal 3/16

OpenStudy (anonymous):

oops! y - 12y^(1/2) + 36 = 0

OpenStudy (anonymous):

Yeah, but it'd still make more sense to just substitute y^2 for x^(3/8)

OpenStudy (anonymous):

ha!

OpenStudy (anonymous):

but then where does substitution come in?